# Brewing Vodka: Distilling Pure Knowledge for Lightweight Threat Detection in Audit Logs

### v1

@@ 4.2. Overall Detection Efficacy Comparsion(Q1)
Table 2 and Table 3 show Vodka ’s performance across various datasets using different teacher models. We observe that Vodka ’s student model consistently improves ACC after distillation from all teacher models. Table 4 compares Vodka with the baselines across multiple datasets. For StreamSpot, we compare Vodka with StreamSpot, ThreaTrace, and MAGIC; for Unicorn Wget, we compare Vodka with Unicorn, Prov-Gem, ThreaTrace, and MAGIC; for DARPA E3 datasets, we compare Vodka with Prov-Gem, ThreaTrace, Log2vec, DeepLog, and MAGIC. Results indicate that Vodka achieves the best or second-best results on all datasets. Specifically, on the StreamSpot, which involves relatively simple attack scenario graphs, Vodka —being an entity-level detector—handles it quite easily. In the Unicorn Wget, which is somewhat more complex compared to StreamSpot but still graph-level, Vodka only slightly lags behind Prov-Gem in Precision (PR), but outperforms all other baselines across all other metrics. In the three DARPA E3, Vodka ranks second only to MAGIC in the Theia and Trace datasets, slightly behind in PR and F1-score.
These results demonstrate that Vodka ’s knowledge distillation approach effectively transfers knowledge from complex GNN teacher models to the student model, enabling the student model to retain high anomaly detection accuracy across multiple detection mechanisms.
Table 4. Vodka ’ comparison results.
@@
Unicorn Wget Vodka 99% 99% 99% 99%
@@
DARPA CADETS Vodka 99% 95% 99% 97%
@@
DARPA TRACE Vodka 98% 95% 99% 97%
@@
DARPA THEIA Vodka 100% 96% 99% 98%
@@ 4.3. Ablation Study (RQ2)
We conducted an ablation study on the CADETS dataset using GCN as the teacher model to investigate the individual contributions of Vodka ’s different submodules, as shown in Figure 4.
@@ 4.4. Hyperparameter Investigation (RQ3)
Training Ratio: The training ratio determines the number of labeled nodes in each batch. To further demonstrate the effectiveness of Vodka, we tested values of 5, 10, 20, 50, 100. The experimental results are shown in Figure 5.(b). As we can observe, the classification performance of the student model improves as the number of labeled nodes increases, reaching near-optimal performance when the batch contains 100 labeled nodes. We attempt to analyze the reasons behind this phenomenon: Increasing the number of labeled nodes introduces more supervised signals, which in turn enhances the efficiency of label inference during PRL. For a semi-supervised system like Vodka, this improvement in label propagation is undoubtedly significant in boosting the model’s performance.
@@ 4.5. Performance Overhead (RQ4)
Vodka aims to perform APTs detection with minimal overhead and flexible deployment in various environments. In real-world deployments, GPUs may be unavailable, and only CPUs may be used. Therefore, we primarily focused on Vodka ’s detection time and memory usage under CPU-only conditions. Table 5 compares Vodka ’s performance overhead with ShadeWatcher and MAGIC on the DARPA E3 TRACE dataset. Results show that Vodka has significant advantages in both detection time and memory usage. At the same training ratio, Vodka is 1.4 times faster than ShadeWatcher and 5.2 times faster than MAGIC.
Given its high detection speed, Vodka could be deployed as a real-time online detector. For example, with the TRACE dataset, which took two weeks to collect, generating 1.37GB of audit logs per day, Vodka only requires 23 seconds per day to complete the APTs detection. This means Vodka ’s lightweight and high-speed detection capabilities make it suitable for daily APTs monitoring in small organizations and enterprises.
@@ 4.6. Robustness Against Adversarial Attacks (RQ5)
As more works focus on provenance-based data analysis, adversarial attacks(Chakraborty et al., 2021; Tramer and Boneh, 2019) designed to evade such provenance-based detectors are also increasing. One common adversarial attack is adversarial mimicry, which alters provenance data and adds benign representations to mimic benign behavior, disguising malicious system entities to evade detection. In this section, we simulate mimicry attacks to evaluate Vodka ’s robustness against adversarial mimicry, using evaluation standards from (Goyal et al., 2023) and (Rehman et al., 2024). Our mimicry attack steps are as follows: first, we select benign system entities from the normal training data. Next, we introduce events and combine them with the benign entities. Finally, we embed the structure of the benign nodes into the attack graph. We compare Vodka (using GCN as the teacher model and CADETS as the dataset) with FLASH(Rehman et al., 2024). Results (Figure 6) show that, as false events increase, FLASH’s anomaly score significantly decreases, indicating its weak robustness against mimicry attacks. In contrast, Vodka maintains an anomaly score above 95% for all levels off false events, demonstrating its superior robustness.
The reasons for this are twofold: first, Vodka is based on entity-level detection, making it highly sensitive to large-scale false event additions. Second, the provenance graph’s neighbor denoising process smooths out some of the strong malicious expressions, making it less susceptible to manipulation.
Table 5. Vodka ’ performance overhead.
@@ 6. Conclusion
In this paper, we innovatively propose a novel APTs detection method called Vodka, which integrates knowledge distillation and provenance-based APTs detection. Vodka introduces a graph Laplacian-based approach for neighbor denoising and signal smoothing on provenance graphs. Additionally, Vodka designs a knowledge distillation framework that distills GNNs knowledge from a complex teacher model into a lightweight student model. The student model combines feature transformation and label propagation, allowing for low-cost detection after training and reconstruction of the attack chain. We evaluate Vodka’s performance on three widely used datasets, demonstrating excellent detection results.

### v2

@@ 4.2. Overall Detection Efficacy Comparsion(Q1)
Table 2 and Table 3 show Winemaking ’s performance across various datasets using different teacher models. We observe that Winemaking ’s student model consistently improves ACC after distillation from all teacher models. Table 4 compares Winemaking with the baselines across multiple datasets. For StreamSpot, we compare Winemaking with StreamSpot, ThreaTrace, and MAGIC; for Unicorn Wget, we compare Winemaking with Unicorn, Prov-Gem, ThreaTrace, and MAGIC; for DARPA E3 datasets, we compare Winemaking with Prov-Gem, ThreaTrace, Log2vec, DeepLog, and MAGIC. Results indicate that Winemaking achieves the best or second-best results on all datasets. Specifically, on the StreamSpot, which involves relatively simple attack scenario graphs, Winemaking —being an entity-level detector—handles it quite easily. In the Unicorn Wget, which is somewhat more complex compared to StreamSpot but still graph-level, Winemaking only slightly lags behind Prov-Gem in Precision (PR), but outperforms all other baselines across all other metrics. In the three DARPA E3, Winemaking ranks second only to MAGIC in the Theia and Trace datasets, slightly behind in PR and F1-score.
These results demonstrate that Winemaking ’s knowledge distillation approach effectively transfers knowledge from complex GNN teacher models to the student model, enabling the student model to retain high anomaly detection accuracy across multiple detection mechanisms.
Table 4. Winemaking ’ comparison results.
@@
Unicorn Wget Winemaking 99% 99% 99% 99%
@@
DARPA CADETS Winemaking 99% 95% 99% 97%
@@
DARPA TRACE Winemaking 98% 95% 99% 97%
@@
DARPA THEIA Winemaking 100% 96% 99% 98%
@@ 4.3. Ablation Study (RQ2)
We conducted an ablation study on the CADETS dataset using GCN as the teacher model to investigate the individual contributions of Winemaking ’s different submodules, as shown in Figure 4.
@@ 4.4. Hyperparameter Investigation (RQ3)
Training Ratio: The training ratio determines the number of labeled nodes in each batch. To further demonstrate the effectiveness of Vodka, we tested values of 5, 10, 20, 50, 100. The experimental results are shown in Figure 5.(b). As we can observe, the classification performance of the student model improves as the number of labeled nodes increases, reaching near-optimal performance when the batch contains 100 labeled nodes. We attempt to analyze the reasons behind this phenomenon: Increasing the number of labeled nodes introduces more supervised signals, which in turn enhances the efficiency of label inference during PRL. For a semi-supervised system like Winemaking, this improvement in label propagation is undoubtedly significant in boosting the model’s performance.
@@ 4.5. Performance Overhead (RQ4)
Winemaking aims to perform APTs detection with minimal overhead and flexible deployment in various environments. In real-world deployments, GPUs may be unavailable, and only CPUs may be used. Therefore, we primarily focused on Winemaking ’s detection time and memory usage under CPU-only conditions. Table 5 compares Winemaking ’s performance overhead with ShadeWatcher and MAGIC on the DARPA E3 TRACE dataset. Results show that Winemaking has significant advantages in both detection time and memory usage. At the same training ratio, Winemaking is 1.4 times faster than ShadeWatcher and 5.2 times faster than MAGIC.
Given its high detection speed, Winemaking could be deployed as a real-time online detector. For example, with the TRACE dataset, which took two weeks to collect, generating 1.37GB of audit logs per day, Winemaking only requires 23 seconds per day to complete the APTs detection. This means Winemaking ’s lightweight and high-speed detection capabilities make it suitable for daily APTs monitoring in small organizations and enterprises.
@@ 4.6. Robustness Against Adversarial Attacks (RQ5)
As more works focus on provenance-based data analysis, adversarial attacks(Chakraborty et al., 2021; Tramer and Boneh, 2019) designed to evade such provenance-based detectors are also increasing. One common adversarial attack is adversarial mimicry, which alters provenance data and adds benign representations to mimic benign behavior, disguising malicious system entities to evade detection. In this section, we simulate mimicry attacks to evaluate Winemaking ’s robustness against adversarial mimicry, using evaluation standards from (Goyal et al., 2023) and (Rehman et al., 2024). Our mimicry attack steps are as follows: first, we select benign system entities from the normal training data. Next, we introduce events and combine them with the benign entities. Finally, we embed the structure of the benign nodes into the attack graph. We compare Winemaking (using GCN as the teacher model and CADETS as the dataset) with FLASH(Rehman et al., 2024). Results (Figure 6) show that, as false events increase, FLASH’s anomaly score significantly decreases, indicating its weak robustness against mimicry attacks. In contrast, Winemaking maintains an anomaly score above 95% for all levels off false events, demonstrating its superior robustness.
The reasons for this are twofold: first, Winemaking is based on entity-level detection, making it highly sensitive to large-scale false event additions. Second, the provenance graph’s neighbor denoising process smooths out some of the strong malicious expressions, making it less susceptible to manipulation.
Table 5. Winemaking ’ performance overhead.
@@ 6. Conclusion
In this paper, we innovatively propose a novel APTs detection method called Winemaking, which integrates knowledge distillation and provenance-based APTs detection. Winemaking introduces a graph Laplacian-based approach for neighbor denoising and signal smoothing on provenance graphs. Additionally, Winemaking designs a knowledge distillation framework that distills GNNs knowledge from a complex teacher model into a lightweight student model. The student model combines feature transformation and label propagation, allowing for low-cost detection after training and reconstruction of the attack chain. We evaluate Winemaking’s performance on three widely used datasets, demonstrating excellent detection results.